Reports 1-1 of 1 Clear search Modify search
MIF (ITF Control)
kiwamu.izumi - 18:19 Tuesday 22 June 2021 (17217) Print this report
AA chassis replaced

Satoshi and Kiwamu,
Related log: 17188
We have swapped an AA unit in the ALS0 rack with a spare, today.

Channel 25 in the old unit showed an anomalously high noise level as reported in 17188. This channel by chance had been used for measuring the phase error for the X PLL. Although we didn't think this is related to the glitch events that we have been trying to hunt, we replaced the unit in order to sort out existing issues.

Resulting noise:
See the screenshot below. The discrepancy between the two channels (15063, Sep. 2020) are now much smaller.  We confirmed that the PLL X guardian can still bring it to the locked state.

A caveat:
Even though the new unit has been thoroughly tested by Shimode san by measuring the transfer functions of all the channels, channel 31 showed a high noise peak at around 1 kHz dominating the rms. The screenshot below shows the spectrum of channel 31 without connecting the signals. The other neighboring channels are also plotted for comparison. Fortunately, this channel is not assigned to any signals. Therefore we will keep using this unit.


Images attached to this report
Comments to this report:
tomotada.akutsu - 0:24 Wednesday 23 June 2021 (17228) Print this report

For my curiosity; your "without connecting the signals" means (1) the input was short to GND or (2) the input was open, which?

takahiro.yamamoto - 2:51 Wednesday 23 June 2021 (17230) Print this report
> channel 31 showed a high noise peak at around 1 kHz
It's no problem.

Ch#31 of ADC#0 is assigned for the timing duotone signal (960+961Hz).
User models cannot assign this channel to other signals.
kiwamu.izumi - 10:26 Wednesday 23 June 2021 (17234) Print this report


I meant the latter, the inputs on the AA filters were left open during the measurement.


Thanks for pointing it out. I completely forgot about the duo-tone.

Search Help